OpenStudy (anonymous):

The following function may be viewed as a composition h(x)=f(g(x)). Find f(x) and g(x). h(x) = sqrt(x^2) - 4 x + 3.

OpenStudy (anonymous):

they want me to find f(x) and g(x)

OpenStudy (anonymous):

i think i need to use the the product rule?

OpenStudy (phi):

Is everything inside the square root? \[ h(x)= \sqrt{x^2-4x+3} \] ?

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

f(x)=sqrt(x) g(x)=x^2-4x+3

OpenStudy (phi):

h(x)= f(g(x)) so if f(x)= sqrt(x) and g(x)= x^2-4x+3 you would "compose" them, by replacing g(x) in f(g(x)) with the definition g(x)= x^2-4x+3 so h(x)= f(x^2-4x+3) that means replace x in f(x)= sqrt(x) with the entire expression x^2-4x+3 to get h(x) = sqrt(x^2-4x+3)

OpenStudy (phi):

btw, you should have written instead of h(x) = sqrt(x^2) - 4 x + 3. h(x) = sqrt(x^2 - 4 x + 3) (with the parens around the whole expression. sqrt(x^2) - 4 x + 3 really means x - 4x+3 = -3x+3 (sqrt(x^2) is x))

OpenStudy (anonymous):

ok so f(X) is is the sqrt of x^2-4x+3?

OpenStudy (lalaly):

no \[f(g(x))=\sqrt{x^2-4x+3}\]

OpenStudy (anonymous):

then what is f(x)?

OpenStudy (lalaly):

\[f(x)=\sqrt{x}\]

OpenStudy (anonymous):

the way u guys found g(X) is that reverse product rule?

OpenStudy (anonymous):

just x or x^2-4x+3?

OpenStudy (phi):

This is function composition. Can you explain what you mean by the product rule?

OpenStudy (anonymous):

oh i mean the chain rule

OpenStudy (phi):

Do you mean h(x) = f(x)*g(x) (where * means multiply)? If so, no, you do not use the product rule. You use the "composition rule"

OpenStudy (anonymous):

the formula is d/d= f(g(x))= f'(g(x))g'(x)

OpenStudy (anonymous):

d/dx

OpenStudy (anonymous):

i havent learned the composition rule yet

OpenStudy (anonymous):

really though just want to find out how to write out f(x) =

OpenStudy (phi):

Say you have 2 functions: f(x)=sqrt(x) g(x)=x^2-4x+3 To compose them to create a new function h(x)= f(g(x)) instead of plugging in a number for x you plug in a function. so f(g(x)) means replace x in f(x) definition with g(x): h(x) = f(g(x)) = sqrt( g(x) ) now replace g(x) with its definition sqrt(g(x) ) = sqrt ( x^2-4x+3) does that make sense?

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

it does makes sense